What velocity would a 60 kg person need to move at in order to have the same momentum as a 600 nm photon?
1
Expert's answer
2020-05-21T13:37:07-0400

Given data

Mass of person is "m=60 kg"

Wavelength of photon is "\lambda=600 nm =600*10^{-9} m"

According to de Broglie hypothesis, the wavelength of matter wave is

"\lambda=\frac{h}{mv}"

The velocity of the person is

"v=\frac{h}{m\lambda}"

"=\frac{6.626*10^{-34}}{(60)(600*10^{-9})}=1.84*10^{-29} m/s"

Hence, the velocity of the person is "v=1.84*10^{-29} m\/s"